OpenStudy (anonymous):

A triangle has sides of and 3. Which could not be the length of the third side if it is a right triangle

OpenStudy (aum):

There should be three numbers for the lengths of three sides. They are missing from the problem.

OpenStudy (anonymous):

well thats goning to be even harder to figure out

OpenStudy (anonymous):

but thanks

OpenStudy (aum):

I am saying you may have mistyped the problem. The first sentence does not make sense: "A triangle has sides of and 3." There should be two other numbers between "of" and "and".

OpenStudy (anonymous):

\[\sqrt{2} \]

OpenStudy (aum):

There should be 3 numbers. Can you post a screenshot of the question?

OpenStudy (anonymous):

is it actually correct tho???

OpenStudy (aum):

It cannot be: \(\large \sqrt{13} \) \(\large ({3})^2 - (\sqrt{2})^2 = (\sqrt{7})^2\). So A is fine. \(\large (\sqrt{2})^2 + (3)^2 = (\sqrt{11})^2 \). So B is fine.

OpenStudy (anonymous):

ok thanks

OpenStudy (aum):

you are welcome.

OpenStudy (anonymous):

so its not 11 because i got 11 wrong

OpenStudy (anonymous):

its 13

OpenStudy (aum):

Oh, you misunderstood my reply above. I was saying the side of the triangle could not be \(\large \sqrt{13}\). So that is the correct answer to the question because it asks for: "Which COULD NOT BE the length of the third side." And I showed how the side of the triangle can be \(\large \sqrt{7}\) and \(\large \sqrt{11}\). So those are not the answers to the question.